OpenStudy (anonymous):

Substitution Property of Equality If y = -5 and 7x + y = 11 , then ______. a. 7(-5) + y = 11 b. 7x - 5 = 11 c. 7x + 5 = 11 d. -5 + y = 11

OpenStudy (anonymous):

Given \frac{ x }{ 5 } + 9 = 11 Prove x = 10 \frac{ x }{ 5 } + 9 = 11 a.____ \frac{ x }{ 5 } = 2 b.____ x = 10 c._____ A. a. Given b. Subtraction Property of Equality c. Division Property of Equality B. a. Given b. Subtraction Property of Equality c. Multiplication Property of Equality C. a. Given b. Addition Property of Equality c. Multiplication Property of Equality D. a. Given b. Addition Property of Equality c. Division Property of Equality

OpenStudy (anonymous):

It is B